Part B: Normal Distribution Probabilities. For all questions, include the probability statement, a sketch, the calculator syn do all of part b please :)
Answer #1

PC Between -0.29 and 1.21) = P(-0.292221.21) - P(241.27)-P (2<-0.29) = P(Z <1.27)-1+ P(Z <0.29) 1.27 = 0.8980-1+ 0.6141 -0.29According to the rules only first four subparts to be answered.